OpenStudy (compassionate):

A bottle of 250 vitamin tablets sells for $3.75. At this rate, what should a bottle of 40 tablets cost? (Choose the answer, marked correctly, from the standard grids.)

OpenStudy (compassionate):

I under the equation looks like this: \[\frac{ 3.75 }{ 250 } = \frac{ x }{ 40 }\] But why is 3.75 over 250?

OpenStudy (anonymous):

Cross multiply (3.75 * 40) 250x= 150; Divide by both sides

OpenStudy (compassionate):

I know how to solve it. I know how to do problems like this. I just never understood why I am putting 250 in the denominator.

OpenStudy (anonymous):

Umm, change it around?

OpenStudy (nincompoop):

oh you did it correctly

OpenStudy (compassionate):

I know I did it correctly... why is 250 in the denominator...

OpenStudy (nincompoop):

let me set another example so that it will be easier

OpenStudy (nincompoop):

it's like saying a pan of pizza with eight slices cost 8 bucks if I were to buy only a slice then we both know intuitively that it'd be a dollar

OpenStudy (nincompoop):

that is 8slices/8bucks = 1slice/1buck

OpenStudy (nincompoop):

let us try another one

OpenStudy (the_fizicx99):

"A bottle of 250 vitamin tablets sells for $3.75" By putting \(\ \dfrac{$3.75}{250}\) you're finding how much each tablet is worth.

OpenStudy (the_fizicx99):

I suspect troll, but flutter it

OpenStudy (compassionate):

Yes... I am aware of how to solve it. I know how to solve these proportions intuitively. That's the problem. I'm asking for insight.

OpenStudy (nincompoop):

I am giving the insight

OpenStudy (nincompoop):

don't listen to fiz

OpenStudy (compassionate):

I appreciate it, Nincompoop. So, we put costs on the top?

OpenStudy (anonymous):

LOL

OpenStudy (the_fizicx99):

xD What.. I explained the reasoning behind putting 3.75/250, not how to solve it..

OpenStudy (compassionate):

Hmm. It kind of wouldn't make sense if we divided 3.75 into 250. Because that would mean each pill would be like 60 dollars.

OpenStudy (nincompoop):

it depends on what you are looking for, comp you put on the numerator that is which you wanted to break into parts, and the number of parts you want to break it into will be the denominator

OpenStudy (anonymous):

http://prntscr.com/4t5bhr

OpenStudy (compassionate):

Melissa went shopping and spent half of her money on shoes, a third on a blouse, a tenth to take her boyfriend to lunch, and she came home with $12. How much did she start out with? So, let me see. x = x/2 + x/3 + x/10 + 12 And we would find the LCM 30 = 15x + 10x + 3x + 360 Now we add like terms, and solve?

OpenStudy (nincompoop):

I want to break a pan of pizza into 8 parts will be interpreted as \( \frac{1pan pizza}{8 slices}\)

OpenStudy (compassionate):

Ah, alright. I get it. I'm just trying to refresh my memory on the basics and stop. Thanks a lot, @nincompoop

OpenStudy (nincompoop):

in your problem, you want to break the total cost into 250 parts to get the cost of each pill

OpenStudy (compassionate):

It's been years since I touched on algebra and arithmetic. :=P

OpenStudy (compassionate):

I understand it Nin. I appreciate your help a lot :)

OpenStudy (compassionate):

If you gave me trig, I could do that easily. I just need to refresh my memory on the basics. Felt I've been neglecting it lately. xc

OpenStudy (nincompoop):

it's cool in chemistry, you will be making good use of fraction and conversion by dimensional analysis

OpenStudy (compassionate):

I made a 93 in chemistry. I don't know how, to be honest. A lot of the stuff is just repetitive for me. It's become just another process when I read something and analyze it. I like to go over these concrete notions I have and try to understand them more.

OpenStudy (jonnyvonny):

@Compassionate I don't know if anyone already said this, but there is no matter in where you put the 250 and the price, so long as you put everything else they same way.

OpenStudy (jonnyvonny):

the same way**

OpenStudy (anonymous):

i have a suggestion

OpenStudy (anonymous):

forget about the ratio for a second compute the price per tablet if 250 cost \(\$3.75\) then each one costs \(\frac{3.75}{250}\)

OpenStudy (anonymous):

that is the cost of one tablet, then 40 tablets cost 40 times as much, \[\frac{40\times 3.75}{250}\]

OpenStudy (anonymous):

and 60 would cost 60 times as much etc

OpenStudy (compassionate):

Shouldn't 40 tablets cost 40 times less? Since we have 250 tablets for 3.75 a piece...

OpenStudy (anonymous):

oh no it is not 250 for $3.75 each!

OpenStudy (anonymous):

it is $3.75 for the whole lot of 250

OpenStudy (compassionate):

Oh, righto right! Hehe! Okay, so I understand now how we can have 250 divided into 3.75 to get the price per-tablet. But what did you do when you multiplied 3.75 by 40 then divided it all by 250?

OpenStudy (compassionate):

If you multiply it by 40, wouldn't that just raise the cost of each tablet?

OpenStudy (anonymous):

\[3.75\div 250\] is the "unit price" i.e. the price per tablet

OpenStudy (anonymous):

you want to know how much whatever number of tablets cost, you multiply it by that number